OpenStudy (anonymous):

if f(x) = 2x then f^{-1} (x)= ?

OpenStudy (anonymous):

it is asking for the inverse, i think? y = 2x inverse is x = 2y x/2 = y x/2 = f^-1(x)

OpenStudy (anonymous):

Let f(x) = y So : y = 2x FInd x from here: \[x = \frac{y}{2}\] Just flip x and y : \[y = \frac{x}{2}\] Here y = \(f^{-1}(x)\)

OpenStudy (anonymous):

point of inverse: x = input, y = output switch x and y (switch the input to the output and vice versa) you will get an "opposite" equation